# A Low-Cost, Low-Power Media Converter Solution for Next-Generation Detector Readout Systems

### v2

@@ 4 Front-End Emulator
A font-end emulator has been developed separately to support the development of next generation back-end electronics, by emulating the output of various detector front-end electronics on FPGA gateware. The emulator is based on the AMD Zynq Ultrscale+ System on Chip, hosted on the ZCU102 evaluation board. The board includes 4 SFP+ cages connected to a Quad of GTH transceivers located on the Zynq chip. Each of the 4 SFP+ ports can be used to emulate one full-speed lpGBT interface. The ZCU102 board also includes an FMC connector which is used to mount a VLDB+ evaluation board, hosting a single lpGBT ASIC chip, which serves as an additional front-end interface.
@@ 6 Achievements and Future Directions
A second prototype is under development. This next iteration will utilize a System on Module (SoM) based on the AMD Zynq UltraScale+, featuring 25Gbps-capable transceivers. The new device is designed to support the processing of up to 48 lpGBT links and output through 5 100GbE uplinks. Multiple of these SoMs could be housed in a 1U rack-mounted enclosure, optimizing space efficiency. This solution aims to further reduce costs while providing a high link density per rack unit.

### v3

@@ 4 Front-End Emulator
A font-end emulator has been developed separately to support the development of next generation back-end electronics, by emulating the output of various detector front-end electronics on FPGA gateware. The emulator is based on the AMD Zynq Ultrascale+ System on Chip, hosted on the ZCU102 evaluation board. The board includes 4 SFP+ cages connected to a Quad of GTH transceivers located on the Zynq chip. Each of the 4 SFP+ ports can be used to emulate one full-speed lpGBT interface. The ZCU102 board also includes an FMC connector which is used to mount a VLDB+ evaluation board, hosting a single lpGBT ASIC chip, which serves as an additional front-end interface.
@@ 6 Achievements and Future Directions
A second prototype is under development. This next iteration will utilize a System on Module (SoM) based on the AMD Zynq UltraScale+, featuring 25Gbps-capable transceivers. The new device is designed to support the processing of up to 48 lpGBT links and output through 5 100GbE uplinks. Multiple of these SoMs could be housed in a 1U rack-mounted enclosure, optimizing space efficiency.
Cost Evaluation
The proposed solution’s cost is compared to the high-end FPGA readout board used in LHCb, the PCIe400, which processes 46 lpGBT links and is priced at \$16,600 [12], resulting in a cost of \$360 per lpGBT link. Most of the PCIe400 cost is attributed to the FPGA, with an estimated \$10,000 for the FPGA and \$6,600 for the rest of the board. Assuming similar costs for optical transceivers222four 12-link Samtec FireFly modules [13] leaves \$4,200 for manufacturing and remaining bill of material costs.
The next iteration of the proposed solution will handle 48 lpGBT links. The selected SoM is the Enclustra XRU90, priced at \$4,123[14]. The SoM absorbs most board costs, and hosting multiple SoMs on the same baseboard enables sharing of expensive components like low-jitter PLLs and voltage regulators. Based on these considerations and the cost of existing SoM baseboards, we estimate an additional \$500 per SoM for baseboard costs.
This analysis demonstrates that the proposed solution is highly cost-effective, with a cost of \$134 per lpGBT link compared to existing designs, as summarized in Table 2.
Board PCIe400 NetGBT
FPGA SoM Cost 10000 4123
Transceivers Cost 2400 2400
PCB Cost 4200 500
Total Board Cost 16600 6423
lpGBT Links 46 48
Cost per lpGBT link 361 134
Table 2: Estimated cost comparison between the PCIe400 FPGA readout board and the NetGBT solution. All prices are in \$USD.